### v1

@@ 4.2.3 Relationship between Edit Actions and Type of Bugs
| Set 1 |
1437 |
1088 |
24.28 |
| Set 2 |
1803 |
1563 |
13.31 |
| Set 3 |
2548 |
1792 |
29.67 |
| Set 4 |
2165 |
1512 |
30.16 |
In short, the developer study demonstrates the real-world benefits of our findings and provides evidence that our findings can enhance the reproducibility of deep learning bugs.
@@ 5 Discussions
In this section, we first provide actionable insights about the reproducibility of deep learning bugs and recommendations for further research (see Section 5.1). We then describe a proof-of-concept tool, dlRepro, that we developed to recommend edit actions and useful information for reproducing deep learning bugs (see Section 5.2).
@@ 5.1 Reproducibility of Deep Learning Bugs
5.2 dlRepro: Proof-of-Concept Tool for Improving Bug Reproduction
From our user study (see RQ3), we observed that the lack of automated debugging tools is one of the most significant challenges in deep learning bug reproducibility. To aid the developers in reproducing deep-learning bugs, we created a proof-of-concept tool, dlRepro 777https://dlrepro.streamlit.app/ (Fig. 2). This tool takes in either a Stack Overflow question ID or a bug report as input. A bug report should include a bug description, code snippets demonstrating the issue, and relevant tags categorizing the bug. Based on the input provided, it recommends the useful information and edit actions needed to reproduce the deep learning bug. We discuss the methodology of the tool below.
Preparing the Bug Reports: Bug reports can be passed to the tool in two ways, which are described below:
Manual Entry: The user can provide the bug description, code snippet, and tags directly to the tool.
Stack Overflow: The user can input the Stack Overflow Question ID into the tool. Using this question ID, our tool utilizes the Stack Overflow API to retrieve the bug description, code snippet, and tags associated with that question.
Tag Analysis: We match the tags provided in the Stack Overflow question to our taxonomy of bug-related tags (Table 1). If there is a perfect match, we determine the bug type directly. However, if there is no match, we use a SentenceTransformer model called all-MiniLM-L6-v2 reimers2019sentence to find the semantic similarity between the tags in the post and our taxonomy tags for different bug types. This allows us to determine the most likely bug type based on the semantic similarity of the tags, even when there is no direct match.
Use case of our POC tool: To demonstrate how our tool can be used to reproduce deep learning bugs, let us consider the Stack Overflow Issue #54028667. This issue 888https://stackoverflow.com/q/54028667 is caused by an incorrect usage of the scikit-learn API. The tags for this issue are ‘python’, ‘scikit-learn’, ‘gridsearch’, ‘gridsearch-cv’. Since none of these tags exactly match the tags from our analysis (Table. 1), the tool computes the semantic similarity to determine the type of bug. Based on semantic similarity values, the tool determines the type of bug to be API. After determining the type of bug, the tool consults our findings from RQ2 (Table. 7 and Table. 8). It presents useful information and key edit actions which can be used to reproduce the bug. The outputs from the tool are shown in Fig. 3 and Fig. 4.
The bug report contained all the information required for reproduction. Hence, by using the recommended edit actions and the complementary code snippets, we could reproduce the bug in 23 minutes, which shows the effectiveness of our findings and the tool. As the bug contained the required information, other people could also reproduce the issue, and the question received an accepted answer within 12 hours. Although our dataset consisted of posts containing ‘tensorflow’, ‘pytorch’, and ‘keras’, we could reproduce a bug related to ‘scikit-learn’. This highlights that our findings and tool are applicable beyond the specific libraries used to build the original dataset. Reproducing the ‘scikit-learn’ bug shows that our approach works even when applied to a library not seen during initial development and testing, highlighting the generalizability of our findings and tool.
Extending and Enhancing Usability of dlRepro: While our study focused on exploring and improving the state of deep learning bug reproducibility, dlRepro was developed as a proof-of-concept tool to demonstrate how our findings can serve as a step towards automated reproduction of deep learning bugs. The tool’s flexible architecture enables continuous offline updates to the underlying algorithm for tag similarities, information category mappings, and edit recommendations without changes to the user-facing interface. It also encourages broader deep-learning developer community contributions. Through this collaborative access model, experts could voluntarily expand the tool’s capabilities by sharing new insights uncovered. Such community participation can lead to gradual, real-world refinements and expansion of both the dlRepro tool and the taxonomy or tags that empower our tool. We have also open-sourced dlRepro’s code to enable direct extensions by the community dlReproCitation , further showcasing the potential for collaborative improvement. While demonstrating the applicability of our analysis, dlRepro establishes a framework for ongoing communal efforts to enhance deep learning bug reproducibility through collective intelligence.
5.3 Comparison with ChatGPT
To assess the effectiveness of our POC tool in deep learning bug reproducibility, we conducted a comparison with ChatGPT, considering the widespread usage of large language models. To do this, we provided ChatGPT with a bug report, supplementary information, and a specific query, as shown in Fig. LABEL:fig:chatGPTQuery. To evaluate ChatGPT’s capability to reproduce deep learning bugs, we supplied it with the identical bug report, code snippets, and other details utilized in our POC tool’s use case (Section 5.2).
The query included an instruction asking ChatGPT for help reproducing the bug, while the query context included the detailed bug description and corresponding code snippet. By providing this additional context, we aimed to enhance ChatGPT’s understanding of the bug. We wanted the ChatGPT to reproduce the bug.
Upon requesting ChatGPT to reproduce the bug, the model provided a potential solution to the problem. Although this solution could potentially resolve the bug, it might not be enough to reproduce the bug, which was our original goal. Unlike ChatGPT’s response, our findings allowed us to reproduce the bug successfully by combining edit actions and useful information, as shown in Section 5.2.
This case study serves as a compelling demonstration of the practicality and value of our findings, even in the era of language models. It highlights how our approach can offer unique insights and solutions that are not readily apparent through language models alone. By leveraging our findings and language models’ capabilities, we can maximize the effectiveness of bug reproducibility techniques and drive advancements in this field.
@@ 6 Threats to Validity
Threats to internal validity pertain to experimental errors and confounding variables during the reproduction of bugs. The possibility of introducing new bugs into the code exists during bug reproduction. We have implemented several strategies and precautions to mitigate these threats in our experimental design. For instance, bug reproduction attempts are carried out in a controlled and isolated environment. We refrain from using hardcoded or fixed values in our code snippets to prevent potential errors during bug reproduction. For instance, during the bug reproduction, we utilize random hyperparameter initialization, construct multiple neural networks, and generate randomized input data that aligns with the original distribution. This methodology mitigates the influence of unforeseen variables on the outcomes of our experiments. Following bug reproduction, we undertake rigorous manual analysis to identify any discrepancies or anomalies in the outcomes. Another threat to validity is the incomplete or incorrect information in the bug report, which may lead to incorrect conclusions about the reproducibility of deep learning bugs. To mitigate this threat, we apply extensive filtration based on the criteria suggested by Moravati et al. defects4ml and Humbatova et al. taxonomyRealFaults , which leads to a clean and generalizable dataset.
@@
Threats to construct validity relate to the use of appropriate metrics for evaluating the results of this user study. We used the reproducibility rate and time to reproduce the deep learning bugs as the metrics to evaluate the benefits of our findings. These metrics directly measure how easily the bugs can be reproduced, which aligns with our goal of making deep learning bugs more reproducible. To ensure these metrics were accurately measured, we had clear criteria for determining whether a bug was successfully reproduced. We also recorded precise timestamps for how long it took participants to reproduce bugs. This helped mitigate threats related to measurement errors or subjective assessments of the metrics. Another threat to the validity might be the accuracy of tags used to categorize the bugs. The incorrect categorization of bugs might lead to incorrect conclusions for different bugs. To mitigate this threat, we only select the tags that are present in the taxonomy and the sub-taxonomies defined by Humbatova et al. taxonomyRealFaults . For example, we specifically chose the “loss-function” as one of the tags to distinguish the training bugs since the Loss Function falls within the sub-taxonomy of Training bugs. Thus, the threats to construct validity were effectively mitigated.
@@ 7 Related Work
Unlike many earlier studies above, we conduct an extensive empirical study to understand the challenges of deep learning bug reproduction and how we can improve the reproducibility of deep learning bugs. We construct a dataset of 668 bugs and manually reproduce 85 bugs, spanning three frameworks, five types of bugs, and 14 architectures. We not only (1) define ten edit actions which can be used to reproduce deep learning bugs but also (2) explore the associations among the type of bugs, the information required to reproduce them, and the specific edit actions that can be used to reproduce them, which makes our work novel. Our findings are also generalizable due to the diversity of our dataset. To ensure transparency and reproducibility, we have made our dataset and replication package publicly available999https://github.com/mehilshah/Bug_Reproducibility_DL_Bugs.
@@ 8 Conclusion & Future Work
The reproducibility of deep learning bugs is a significant challenge for software practitioners, given that these bugs can lead to severe consequences. In this paper, we conduct an empirical study by reproducing 85 deep learning bugs and identifying ten key edit actions that could reproduce these bugs. Furthermore, we investigate the relationships among bug types, information categories, and edit actions. To validate our findings, we conducted a user study with 22 participants. The developers equipped with our recommended edit actions and information could reproduce 22.92% more bugs using 24.35% less time compared to the developers without such hints. This demonstrates the real-world value of our findings in improving the reproducibility of deep learning bugs. Finally, we provide actionable insights about the state of deep learning bugs that could improve the reproducibility of deep learning bugs. To demonstrate the practical use of our findings, we built a proof-of-concept tool, dlRepro, that takes bug reports and recommends the edit actions and useful information for deep learning bugs. Moreover, we also conduct a comparison with ChatGPT, which fails to reproduce the bug, while our methodology succeeds. This demonstrates the real-world value of our findings in improving the reproducibility of the deep learning bugs. For future work, we plan to expand our study to additional deep learning frameworks and model types. We will also improve dlRepro with more advanced natural language processing capabilities to extract key information from bug reports. Finally, we aim to integrate dlRepro into developer tools and workflows to provide automated reproducibility support during deep learning development. Furthermore, our findings open up the ability to create automated systems for reproducing deep learning bugs. By training machine learning models on our dataset of reproduction actions matched to bug reports, we can build AI agents that take bug reports as input and automatically generate reproducibility scripts, similar to the work by White et al. reproducibleBugReports . This would significantly reduce the manual effort required to reproduce bugs and enable large-scale debugging and testing of deep learning systems.

### v2

@@ 5.2.3 Relationship between Edit Actions and Type of Bugs
| Set 1 | 1437 | 1088 | 24.28 |
| Set 2 | 1803 | 1563 | 13.31 |
| Set 3 | 2548 | 1792 | 29.67 |
| Set 4 | 2165 | 1512 | 30.16 |
5.3.1 Impact of Hints on Bug Reproducibility
To determine the impact of hints, we constructed a generalized linear model. This allowed us to test the statistical significance of multiple independent factors on bug reproducibility (i.e., dependent variable). The factors were experience with deep learning bug fixing, experience with deep learning, profession, and the presence/absence of hints.
| Variable | Estimate | Std. Error | t value | Pr t |
| Intercept | 0.4734 | 0.104 | 4.565 | 0.000 |
| Hints | 0.4011 | 0.086 | 4.651 | 0.000 |
| Deep Learning Bug Fixing Experience | 0.1085 | 0.142 | 0.763 | 0.445 |
| Deep Learning Experience | 0.0831 | 0.165 | 0.504 | 0.614 |
| Profession (Academia/Industry) | -0.2780 | 0.160 | -1.737 | 0.082 |
Based on the regression results in Table 12, we can see that the presence of hints had a statistically significant positive effect on the reproducibility of deep learning bugs (p 0.001).
On the other hand, factors like experience with deep learning bug fixing, general deep learning experience, and profession (academia vs industry) did not significantly influence reproducibility. However, the participants’ profession (academia or industry) demonstrates a marginal negative effect on the reproducibility of deep learning bugs. (p = 0.082).
Overall, these results highlight that the presence of targeted hints has a significant positive impact on the reproducibility of deep learning bugs.
5.3.2 Impact of Developer Experience on Deep Learning Bug Reproducibility
To perform co-factor analysis, we construct another GLM model. We use deep-learning bug-fixing experience as a factor to evaluate its impact on the reproducibility of deep-learning bugs. The results of the analysis are shown in Table. 13.
| Variable | Coef | Std. Err | z value | Pz |
| Intercept | 0.7839 | 0.300 | 2.610 | 0.009 |
| DLBugFixExp_0[T.True] | -0.2730 | 0.597 | -0.457 | 0.648 |
| DLBugFixExp_1[T.True] | 0.7422 | 0.460 | 1.612 | 0.107 |
| DLBugFixExp_2[T.True] | 0.3147 | 0.651 | 0.484 | 0.629 |
The results from our generalised linear model suggest that years of experience in deep learning bug fixing do not have a statistically significant effect on the reproducibility of deep learning bugs. In particular, both developers with 1-5 years of experience and those with 5-10 years of experience achieved slightly higher bug reproducibility than those with less than 1 years of experience. However, neither of these experience levels have a statistically significant impact on bug reproducibility according to the model (p 0.05). On the other hand, as shown in Table. 12, the presence of targeted hints was found to have a statistically significant positive effect on bug reproducibility (p 0.001). Taken together, the above results indicate that the availability of specific guidance, rather than years of experience, is the major factor that can help the developers reproduce deep-learning bugs.
Summary of RQ3: In the user study, developers were assigned to either the experimental or control group where the experimental group received our suggested edit actions and component information. The experimental group reproduced 22.92% more deep learning bugs and decreased reproduction time by 24.35% compared to the control group. This demonstrates that the identified edit actions and component information substantially improve the reproducibility rate and reduce the time needed to reproduce deep learning bugs.
@@ 6 Discussions
In this section, we first provide actionable insights about the reproducibility of deep learning bugs and recommend potential directions for further research (see Section 6.1). We then demonstrate how our findings can improve the reproducibility of the DL bugs with the use of large language models (e.g., Llama 3) (see Section 6.2).
@@ 6.1 Reproducibility of Deep Learning Bugs
6.2 Improving the Reproducibility of Deep Learning Bugs using LLMs
In this section, we further aim to verify the practical utility of our findings in bug reproduction. To achieve this, we leverage the capabilities of a large language model (LLM) and offer more actionable recommendations to developers.
We conduct experiments using the LLaMA3, a 70B parameter LLM, to evaluate its effectiveness in assisting developers with the reproduction of deep learning bugs. To conduct these experiments, we construct a dataset of 40 bugs. We used stratified random sampling and collect 40 of our manually reproduced bugs, with the following distribution: 10 Training, 10 API, 10 Tensor and 10 Model Bugs. The prompt for the LLaMA3 model contained four key inputs: the bug report describing the issue, relevant code snippets providing context, the instruction to generate a code snippet that could reproduce the reported bug, and, crucially, the edit actions and component information derived from our findings. The prompt for the experiment can be found in our replication package replicationPackage . Once the model generates the code snippets, we check their ability to reproduce the bugs through manual execution and comparison with the erroneous behaviour outlined in the bug reports. Furthermore, we compare the generated code with the ground truth code snippets curated during our manual bug reproduction process.
The results of our experiments are as follows:
Performance without Augmentation: When the LLM was provided with only the bug report and code snippet, its suggestions were less useful. Its generated code snippets lack correctness, despite being compilable and executable. The absence of guidance from our findings diminishes the LLM’s performance, with the model generating reproducible code snippets for only 25 out of the 40 bugs (62.5%).
Performance with Augmentation using Hints: When guided by our edit actions and component information, the LLM generates reproducible code snippets for 33 out of the 40 bugs (82.5%). Analysis of the generated code snippets reveals that the model effectively understands the bug, uses appropriate edit actions, and generates complete, compilable, and executable code snippets. Additionally, in some scenarios, the model provides a textual description of the circumstances under which the bug might be reproducible, offering valuable extra information to developers. Code snippets generated by Llama3, which demonstrate the model’s improved understanding after augmentation using hints, are also available in the replication package replicationPackage .
Manual Analysis of Generated Code Snippets: We perform manual analysis of the generated code snippets to understand the behaviour of LLMs. From the manual analysis of the generated code snippets, we observed that the LLM performs relatively well in generating code snippets for Training and Model bugs, even without augmentation. This behaviour could be explained by the high prevalence of these bugs in real-life scenarios taxonomyRealFaults . On the other hand, the LLM faces more challenges in generating the code snippets for reproducing the Tensor and API bugs without guidance. Tensor bugs usually involve complex manipulations of tensor shapes and dimensions, which might not be fully understood by LLMs. Similarly, API bugs are sensitive to constant updates to the API documentation. The frequent changes to the API documentation make it harder for LLMs to stay up-to-date and generate accurate code snippets to reproduce these bugs. However, when provided with our guidance, the reproducibility of API bugs goes up from 56% to 76%, and Tensor bugs reproducibility goes up from 58% to 75%. This demonstrates the utility of our findings in improving the ability of language models to generate code snippets for bugs that are relatively difficult to reproduce.
Our experiments demonstrate that augmenting the LLM with our suggested edit actions and component information enhances its effectiveness in reproducing deep learning bugs by 20%. This improvement underscores the potential of integrating our findings with LLMs to provide developers with more actionable guidance for reproducing deep learning bugs.
6.2.1 Statistical Significance Tests for our Experiments
To assess the statistical significance of the LLM results, we employ McNemar’s test pembury2020effective , which is designed to detect differences between paired samples of two binary variables. The null hypothesis of McNemar’s test is that the LLM’s bug reproduction ability is the same with or without hints. We organize the samples in a 2x2 contingency table, as shown in Tab. 14. We explain the values present in the table in detail below.
23: This value indicates the number of bugs that were reproduced with/without the hints.
2: This value indicates the number of bugs that were reproduced in the absence of the hints, but could not be reproduced in the presence of the hints.
10: This value indicates the number of bugs that were reproduced in the presence of the hints, but could not be reproduced in the absence of the hints.
5: This value indicates the number of bugs that could not be reproduced with/without the hints.
| Using Hints: | Using Hints: | Total | |
| Reproducible | Not Reproducible | ||
| Without Hints: Reproducible | 23 | 2 | 25 |
| Without Hints: Not Reproducible | 10 | 5 | 15 |
| Total | 33 | 7 | 40 |
It is important to note that the LLM may struggle to reproduce some bugs without full code context, even with our augmentation. Despite these limitations, our results, when compared to the LLM’s performance without augmentation, yield a McNemar’s test p-value of 0.043, which is less than 0.05. The results provide strong evidence to reject the null hypothesis and accept the alternate hypothesis, indicating that the LLM has a statistically significant improved ability to reproduce bugs when provided with hints compared to without hints.
The recommendations of LLMs are stochastic by design, and different runs with the same prompt can produce different reproducible snippets. This introduces an additional source of variability which may invalidate our previous observations. To address this issue, we repeat the code generation task for each bug five times and record their rates of success in bug reproduction. Since the bug reproduction rates are continuous percentage values rather than binary outcomes, the McNemar test cannot be used. To overcome this limitation, we use the RM-ANOVA test rutherford2011anova . RM-ANOVA is suitable for this scenario because we have a within-subjects design with one independent variable (augmentation with hints or not) and one dependent variable (bug reproduction rate) measured repeatedly across trials. The test yields a significant p-value of , further demonstrating the positive impact of hints on bug reproduction abilities of large language models.
@@ 7 Threats to Validity
Threats to internal validity pertain to experimental errors and confounding variables during the reproduction of bugs. The possibility of introducing new bugs into the code exists during bug reproduction. We have implemented several strategies and precautions to mitigate these threats in our experimental design. For instance, bug reproduction attempts are carried out in controlled and isolated Python virtual environments. We refrain from using hardcoded or fixed values in our code snippets to prevent potential errors during bug reproduction. For instance, during the bug reproduction, we utilize random hyperparameter initialization, construct multiple neural networks, and generate randomized input data that aligns with the original distribution. This methodology mitigates the influence of unforeseen variables on the outcomes of our experiments. To further mitigate this threat, we ran the updated code snippets five times for all the edit actions. Following bug reproduction, we undertake rigorous manual analysis to identify any discrepancies or anomalies in the outcomes.
Another threat to validity is the incomplete or incorrect information in the bug report, which could impact the applicability of certain edit actions suggested by our technique. In particular, edit actions that require additional context from the report, such as those involving input data generation, neural network construction, dataset procurement, and downloading models/tokenizers, may be limited by missing details in the reports. However, the majority of our suggested edit actions focus on code modifications that can be derived from the provided code snippets without relying heavily on external context. Therefore, while missing information may constrain our capability for data and model-driven actions, our core technique and findings remain widely applicable for many code-focused edit recommendations. To mitigate this threat further, we apply extensive filtration based on the criteria suggested by Moravati et al. defects4ml and Humbatova et al. taxonomyRealFaults , which leads to a clean and generalizable dataset.
Finally, the way in which the contextual information is presented to the developers could potentially threaten the validity of the control and experiment groups in the user study. To mitigate this threat, we ensure that the contextual information is presented in an identical format to the control and experiment groups.
Threats to construct validity relate to the use of appropriate metrics for evaluating the results of a user study. In the user study, we used the reproducibility rate and time to reproduce the deep learning bugs as the metrics to evaluate the benefits of our findings. To measure these accurately, we had participants justify why and how they used the edit actions or found certain information important. We verified these justifications against our ground truth to determine if the bug was successfully reproduced. After checking the user justifications, we calculated the reproducibility rate and spent time. By including user justifications and verifying them, we ensured that the metrics directly measured how easily bugs could be reproduced. We also had clear criteria for determining reproduction success and used Opinio for precise timing data Opinio , mitigating threats related to measurement errors or subjective assessments.
@@
Another threat to the validity might be the accuracy of tags used to categorize the bugs. The incorrect categorization of bugs might lead to incorrect conclusions for different bugs. To mitigate this threat, we only select the tags that are present in the taxonomy and the sub-taxonomies defined by Humbatova et al. taxonomyRealFaults . For example, we specifically chose the “loss-function” as one of the tags to distinguish the training bugs since the Loss Function falls within the sub-taxonomy of Training bugs. Thus, the threats to construct validity were effectively mitigated.
@@ 8 Related Work
Unlike many earlier studies above, we conduct an extensive empirical study to understand the challenges of deep learning bug reproduction and how we can improve the reproducibility of deep learning bugs. We construct a dataset of 668 bugs and manually reproduce 148 bugs, spanning three frameworks, five types of bugs, and 14 architectures. We not only (1) define ten edit actions which can be used to reproduce deep learning bugs but also (2) explore the associations among the type of bugs, the information required to reproduce them, and the specific edit actions that can be used to reproduce them, which makes our work novel. Our findings are also generalizable to Tensorflow and PyTorch bugs due to the diversity of our dataset. To ensure transparency and reproducibility, we have made our dataset and replication package publicly available777https://github.com/mehilshah/Bug_Reproducibility_DL_Bugs.
@@ 9 Conclusion & Future Work
The reproducibility of deep learning bugs is a significant challenge for software practitioners, given that these bugs can lead to severe consequences. In this paper, we conduct an empirical study by reproducing 148 deep learning bugs and identifying ten key edit actions that could reproduce these bugs. Furthermore, we investigate the relationships among bug types, component information, and edit actions. To validate our findings, we conducted a user study with 22 participants. The developers equipped with our recommended edit actions and information could reproduce 22.92% more bugs using 24.35% less time compared to the developers without such hints. This demonstrates the real-world value of our findings in improving the reproducibility of deep learning bugs. Finally, we provide actionable insights about the state of deep learning bugs that could improve the reproducibility of deep learning bugs. To demonstrate the practical use of our findings, we show how they aid large language models in improving the reproducibility of deep-learning bugs. For future work, we plan to expand our study to additional deep-learning frameworks and model types. Moreover, we plan to conduct a large-scale study to determine the sufficiency of information in deep learning bug reports, similar to the study conducted by Mondal et al. mondal2024can . Using these insights, we will build a tool that automatically completes the reports for better bug reproducibility. Furthermore, our findings open up the ability to create automated systems for reproducing deep learning bugs. By training machine learning models on our dataset of reproduction actions matched to bug reports, we can build AI agents that take bug reports as input and automatically generate reproducibility scripts, similar to the work by White et al. reproducibleBugReports . This would significantly reduce the manual effort required to reproduce bugs and enable large-scale debugging and testing of deep learning systems.
@@
Acknowledgements
We would like to thank Usmi Mukherjee for her invaluable contributions during the manual bug reproduction process as an independent collaborator.
Appendix A Flowchart for Manual Analysis
The flowchart in Fig. 5 illustrates a systematic process for analyzing Stack Overflow posts related to deep learning bugs. It begins by determining if the post discusses a deep-learning bug. If it does, the bug category is identified based on the symptoms, potential root causes, post tags and explanation by the user. The flowchart then guides the user through a series of checks to assess the reproducibility of the bug, considering factors such as the completeness of the code snippet, availability of necessary data, presence of relevant logs or error messages, specification of hyperparameters, and clarity of the model structure. We then try to reproduce the bug. If the reproduction is successful with extra steps, the steps, information, and assumptions needed are documented, or else the bug is marked as non-reproducible.